Find the present value of an ordinary annuity for;



(a) P5500 per quarter for 6.5years at 5.6% compounded quarterly.



(b)P150 per month for 3years at 8% compounded semi-annually.


1
Expert's answer
2021-11-22T11:06:07-0500

a)


"5500=S\times\frac{1.4(1+1.4)^{26}}{(1+1.4)^{26}-1}"

"S=3928.57"

b)


"150=S\times\frac{1.3(1+1.3)^{6}}{(1+1.3)^{6}-1}"

"S=115.38"